OpenStudy (anonymous):

Which of the following is the slope between the points (-2,4) and (2,-2)?

OpenStudy (anonymous):

Use this equation \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] where y2=-2, y1=4, x2=2 and x1=-2 What did you get?

OpenStudy (anonymous):

This equation you always use to find the slope of a line given 2 points

OpenStudy (anonymous):

-2/3?

OpenStudy (anonymous):

or 2/3?

OpenStudy (anonymous):

Nope it is -3/2 (-2-4) = -6 2-(-2) = 2+2 = 4 -6/4 simplified is -3/2

OpenStudy (anonymous):

what about (0,-7) and (4,12)

OpenStudy (anonymous):

y2=12, y1=-7, x2=4,x1=0 y2-y1 = (12-(-7)) = 12+7 = 19 x2-x1=4-0 = 4 m = 19/4

OpenStudy (anonymous):

Do u have any other questions?

OpenStudy (anonymous):

Identify the correct x intercept and the y intercept of the equation x+7y=-21

OpenStudy (anonymous):

To find the y-intercept let x=0 and solve for y To find the x-intercept let y=0 and solve for x What did you get when u applied the principle above?

OpenStudy (anonymous):

I am like so confused haha im terrible at math

OpenStudy (anonymous):

This is what I meant: So you have x+7y=-21 Find y-intercept by letting x=0: 0+7y=-21, then solve for y Find x-intercept by letting y=0: x+7(0)=-21, then solve for x

OpenStudy (anonymous):

so x intercept would be (21,0) and y would be (0,3)

OpenStudy (anonymous):

You forgot your negative signs: x-intercept: (-21,0) y-intercept: (0,-3)

OpenStudy (anonymous):

Do you have any more questions?